Wednesday, 19 August

Special Activities

\underline{14:00 – 18:00 H 104}

Mathematical Software Chair: Winfried Neun, ZIB, Berlin, Germany In this session, a broad spectrum of mathematical software including symbolic and numerical systems as well as interfaces and problem solutions will be presented. Details of the presentations can be found on the ICM’98 web server at URL http://elib.zib.de/ICM98/MathSoftware .

14:00 – 14:30 John J. Cannon et. al., University of Sydney, Australia MAGMA V 2.3

14:40 – 15:10 Gert-Martin Greuel, Gerhard Pfister et. al., Universität Kaiserslautern, Germany Singular

15:15 – 15:45 Lorenzo Robbiano et. al., University of Genova, Italy CoCoA

15:45 – 16:15 Break

16:15 – 16:45 Michael Trott, Wolfram Research, Inc., Illinois, USA Solving Problems with Mathematica

16:50 – 17:20 Ulrich Nowak, Uwe Pöhle et. al., ZIB Berlin, Germany ODELab

17:30 – 18:00 Andreas Dolzmann and Thomas Sturm, Universität Passau, Germany REDLOG

ICM’98 – Program  1